OpenStudy (anonymous):

Determine whether the function has an inverse function. : y=lnx^2

OpenStudy (anonymous):

\[\ln(x^2)~~\text{or}~~(\ln x)^2~~?\]

OpenStudy (anonymous):

It doesnt have any perenthesis

OpenStudy (anonymous):

If that's the case, I'd assume the first one.

OpenStudy (anonymous):



OpenStudy (anonymous):

The usual procedure is to flip the variables, then solve for \(y\). \[y=\ln(x^2)~~\Rightarrow~~x=\ln(y^2)\]

OpenStudy (anonymous):

got that

OpenStudy (anonymous):

but how fo you solve for y?

OpenStudy (anonymous):

\[x=\ln(y^2)\\ x=2\ln y\\ \frac{x}{2}=\ln y\\ y=\cdots \]

OpenStudy (anonymous):

In my book, the answer says there is no inverse... Im assuming that's because if you graph the function it doesn't pass the horizontal/one-to-one test But how would I know that by solving it algebraically like that I have no idea how to graph an equation like that either

zepdrix (zepdrix):

https://www.desmos.com/calculator/fw7vmht6fc Oh I see what you mean. Look at the graph I provided. Our red function, when reflected over y=x (in green) gives us the inverse function in purple. See how we lost one of the branches? :O (Otherwise it wouldn't pass the horizontal line test, as you said.) Algebraically? Hmmm..